On January 1, 2021, the company obtained a $3 million loan with a 11% interest rate. The building was completed on September 30, 2022. Expenditures on the project were as follows: January 1, 2021 $ 1,290,000 March 1, 2021 780,000 June 30, 2021 240,000 October 1, 2021 690,000 January 31, 2022 1,080,000 April 30, 2022 1,395,000 August 31, 2022 2,520,000 On January 1, 2021, the company obtained a $3 million construction loan with a 11% interest rate. Assume the $3 million loan is not specifically tied to construction of the building. The loan was outstanding all of 2021 and 2022. The company’s other interest-bearing debt included two long-term notes of $5,800,000 and $7,800,000 with interest rates of 6% and 8%, respectively. Both notes were outstanding during all of 2021 and 2022. Interest is paid annually on all debt. The company’s fiscal year-end is December 31. Required: 1. Calculate the amount of interest that Mason should capitalize in 2021 and 2022 using the weighted-average method. 2. What is the total cost of the building? 3. Calculate the amount of interest expense that will appear in the 2021 and 2022 income statements.
Answer #1
2021 2022
Interest Capitalized $              1,75,028 $   3,05,174
Interest Expense $            11,26,972 $   9,96,826
Total cost of building $            84,75,202
Workings:
Weighted Average rate of all debt:-
$            30,00,000 X 11% = $          3,30,000
$            58,00,000 X 6% = $          3,48,000
$            78,00,000 X 8% = $          6,24,000
$        1,66,00,000 $       13,02,000
Weighted Average rate of all debt = 7.84%
($1302000/ $16600000)
Expenditure for 2021
Jan 1,2021 $            12,90,000 X 12/12 = $       12,90,000
March 1, 2021 $              7,80,000 X 10/12 = $          6,50,000
June 30, 2021 $              2,40,000 X 6/12 = $          1,20,000
October 1, 2021 $              6,90,000 X 3/12 = $          1,72,500
$            30,00,000 $       22,32,500
Interest Capitalized in 2021
$            22,32,500 X 7.84% = $          1,75,028
Expenditure for 2022
Jan 1, 2022 ($3000000 + $175028) $            31,75,028 X 9/9 = $       31,75,028
Jan 31, 2022 $            10,80,000 X 8/9 = $          9,60,000
April 30, 2022 $            13,95,000 X 5/9 = $          7,75,000
August 31. 2022 $            25,20,000 X 1/9 = $          2,80,000
$            81,70,028 $       51,90,028
Interest Capitalized in 2022
$                      51,90,028 X 7.84% X 9/12 = $          3,05,174
Cost of Building
Expenditure for 2021 = $       30,00,000
Interest Capitalized in 2021 = $          1,75,028
Expenditure for 2022 = $       49,95,000
Interest Capitalized in 2022 = $          3,05,174
Total cost of building = $       84,75,202
Interest Expense for 2021:
Total Interest Incurred = $       13,02,000
Less : Interest Capitalized = $          1,75,028
2021 Expense = $       11,26,972
Interest Expense for 2022:
Total Interest Incurred = $       13,02,000
Less : Interest Capitalized = $          3,05,174
2022 Expense = $          9,96,826